under the same assumptions underlying the model in (1), determine a differential equation for the population P(t) of a country when individuals are allowed to immigrate into the country at a constant rate r>0. What is the differential equation for the population P(t) of the country when individuals are allowed to emigrate from the country at a constant rate r>0?

1
Expert's answer
2021-11-11T18:44:21-0500

When there is no immigration and emigration in a country, the differential equation of growth of population, P(t) satisfies the model

"\frac{dP}{dt}=kP," for some constant k > 0.


(I) When individuals are allowed to immigrate at constant rate r > 0, the differential equation will be

"\frac{dP}{dt}=kP + r"


(ii) When individuals are allowed to emigrate at constant rate r > 0, the differential equation will be

"\frac{dP}{dt}=kP - r"